# GSFF-SLAM: 3D Semantic Gaussian Splatting SLAM via Feature Field

### v1

@@ 3.2 Tracking and Mapping Optimization
$\frac{d\mu_{c}}{dT_{CW}}=[\textit{{I}},-\mu^{\times}],\frac{dW}{dT_{CW}}=[\textit{{0}},-W^{\times}],$ (4)
@@ 4.2 Evaluation of SLAM Metrics
SplaTAM [12] 0.53 0.23 0.36 0.43 0.64 0.33 0.44 0.35 0.413
MonoGS [13] 0.45 0.37 0.27 0.26 0.97 0.39 0.34 0.26 0.414
SGS-SLAM [21] 0.44 0.41 0.52 0.82 0.73 0.32 2.55 1.30 0.885
GSFF-SLAM(Ours) 0.49 0.40 0.32 0.23 0.81 0.31 0.33 0.23 0.390
Table 1: Comparison of tracking performance on Replica dataset [22]. We report ATE RMSE[cm]$\downarrow$ for different sequences. The best results are highlighted as first, second, and third.
@@
NICE-SLAM [9] 12.0 14.0 10.9 13.4 6.2 11.3
Vox-Fusion [10] 16.6 24.2 27.3 23.3 9.4 20.2
Point-SLAM [11] 10.2 7.8 22.2 14.8 9.5 12.9
DNS-SLAM [19] 12.1 5.5 35.6 10.0 6.4 13.9
SplaTAM [12] 12.8 14.3 15.3 13.1 7.8 12.6
MonoGS [13] 9.8 6.4 10.7 23.8 8.1 11.8
SGS-SLAM [21] 12.6 14.0 17.6 13.0 8.0 13.0
@@
GSFF-SLAM(Ours) 17.0 7.8 9.6 21.9 5.6 12.4
@@
GSFF-SLAM(Ours) 1.61 4.98 7.29 1.61 1.96 3.73
@@
Table 4: Quantitative comparison of rendering performance on the Replica [22], ScanNet [46], and TUM-RGBD [47] datasets.
Tracking Accuracy. On the Replica dataset, we evaluate the tracking performance of our method compared to the state-of-the-art methods. As shown in Table 4.2, our method achieves the best tracking accuracy with an average ATE RMSE of 0.390 cm. In the high-quality synthetic environments with accurate depth information and minimal motion blur, our method achieves submillimeter tracking accuracy that is comparable to other state-of-the-art 3DGS-based methods, while consistently outperforming NeRF-based methods. Compared to the NeRF-based SLAM, our method reduces the tracking error by more than 25% from 0.525cm to 0.390cm. Compared to the recent SplaTAM and MonoGS, which achieve average ATE RMSE of 0.413cm and 0.414cm respectively, our method reduces the tracking error by 5.6% and 5.8%. On real-world datasets, Table 2 and Table 3 present the quantitative comparison of tracking accuracy. Our method achieves performance with average ATE RMSE of 12.4cm on ScanNet and 3.73cm on TUM-RGBD. In the ScanNet dataset, we observe certain limitations of our approach. Specifically, in scene0000 ($>$5000 frames) and scene0181, the tracking performance exhibits significant degradation. For scene0000, which represents the longest sequence in our evaluation, we observe gradual drift in trajectory estimation after approximately 3000 frames, primarily due to the absence of loop closure detection. Meanwhile, scene0181 presents additional challenges due to significant motion blur, which adversely affects the mapping quality. These observations highlight potential areas for future improvement, particularly in handling extended sequences and scenes with degraded image quality. On the TUM-RGBD dataset, we conduct comprehensive experiments to further verify the robustness of our approach. Our method achieves competitive performance with an average ATE RMSE of 3.49cm, ranking second among all learning-based methods. It is noteworthy that our algorithm performs poorly in the scene with motion blur, fr1/desk2, where the ATE RMSE of our method is 7.30 cm. In comparison, the NeRF-based SNI-SLAM [20] and the 3DGS-based LoopSplat [48] with loop closure detection achieve ATE RMSE values of 4.35 cm and 3.54 cm, respectively.

### v2

@@ 3.2 Tracking and Mapping Optimization
$\frac{d\mu_{c}}{dT_{CW}}=[\textit{{I}},-\mu_{c}^{\times}],\frac{dW}{dT_{CW}}=[\textit{{0}},-W^{\times}],$ (4)
@@ 4.2 Evaluation of SLAM Metrics
SplaTAM [12] 0.47 0.27 0.29 0.32 0.72 0.31 0.40 0.29 0.384
MonoGS [13] 0.36 0.19 0.25 0.12 0.81 0.33 0.22 0.29 0.321
SGS-SLAM [21] 0.44 0.41 0.52 0.46 0.43 0.32 0.39 0.37 0.418
GSFF-SLAM(Ours) 0.35 0.24 0.25 0.15 0.54 0.34 0.29 0.33 0.311
Table 1: Comparison of tracking performance on Replica dataset [22]. We report ATE RMSE[cm]$\downarrow$ for different sequences.
@@
NICE-SLAM [9] 12.0 14.0 10.9 13.4 6.2 11.30
Vox-Fusion [10] 16.6 24.2 27.3 23.3 9.4 20.14
Point-SLAM [11] 10.2 7.8 22.2 14.8 9.5 12.90
DNS-SLAM [19] 12.1 5.5 35.6 10.0 6.4 13.92
SNI-SLAM [20] 6.9 7.4 - - 4.7 -
SplaTAM [12] 12.8 10.1 12.1 11.1 7.5 10.72
MonoGS [13] 9.8 6.4 10.7 23.8 8.1 11.76
SGS-SLAM [21] 12.6 14.0 17.6 13.0 8.0 13.04
@@
GSFF-SLAM(Ours) 9.6 7.8 9.6 21.9 5.6 10.90
@@
GSFF-SLAM(Ours) 1.61 7.30 7.29 1.61 1.96 3.95
@@
Table 4: Quantitative comparison of rendering performance on the Replica [22], ScanNet [46], and TUM-RGBD [47] datasets. The best results are highlighted as first, second, and third.
Tracking Accuracy. On the Replica dataset, we evaluate the tracking performance of our method compared to the state-of-the-art methods. As shown in Table 4.2, our method achieves the best tracking accuracy with an average ATE RMSE of 0.311cm. In the high-quality synthetic environments with accurate depth information and minimal motion blur, our method achieves submillimeter tracking accuracy that is comparable to other state-of-the-art 3DGS-based methods, while consistently outperforming NeRF-based methods. Compared to the NeRF-based SLAM, our method reduces the tracking error by more than 25% from 0.456cm to 0.311cm. Compared to the recent SplaTAM and MonoGS, which achieve average ATE RMSE of 0.384cm and 0.321cm respectively, our method reduces the tracking error by 19.0% and 3.1%. On real-world datasets, Table 2 and Table 3 present the quantitative comparison of tracking accuracy. Our method achieves performance with average ATE RMSE of 10.90cm on ScanNet and 3.95cm on TUM-RGBD. In the ScanNet dataset, we observe certain limitations of our approach. Specifically, in scene0000 and scene0181, the tracking performance exhibits significant degradation. For scene0000, which represents the longest sequence in our evaluation, we observe gradual drift in trajectory estimation after approximately 3000 frames, primarily due to the absence of loop closure detection. Meanwhile, scene0181 presents additional challenges due to significant motion blur, which adversely affects the mapping quality. These observations highlight potential areas for future improvement, particularly in handling extended sequences and scenes with degraded image quality. On the TUM-RGBD dataset, we conduct comprehensive experiments to further verify the robustness of our approach. Our method achieves competitive performance with an average ATE RMSE of 3.95cm, ranking among all learning-based methods. It is noteworthy that our algorithm performs poorly in the scene with motion blur, fr1/desk2, where the ATE RMSE of our method is 7.30 cm. In comparison, the NeRF-based SNI-SLAM [20] achieve ATE RMSE values of 4.35 cm, respectively.